# The Future Outcome Reasoning and Confidence Assessment Benchmark

### v2

@@ 6 Related Work
Recent forecasting benchmarks focus on event prediction but largely overlook confidence calibration. OpenForecast (Wang et al., 2025) introduces a large-scale dataset for open-ended, multi-step event forecasting but does not assess model confidence. ForecastBench (Karger et al., 2024) evaluates binary (Yes/No) forecasting by prompting models for direct probability estimates, but since it queries each option separately, the assigned probabilities do not necessarily sum to 1, leading to potential inconsistencies. Neither benchmark systematically evaluates confidence calibration, a crucial aspect for reliable forecasting in real-world applications.
Beyond forecasting, several benchmarks assess language models’ reasoning and inference capabilities. COPA (Roemmele et al., 2011) evaluates causal reasoning by presenting a premise and two alternatives, requiring models to select the more plausible cause or effect. HellaSwag Zellers et al. (2019) challenges models with sentence completion tasks that demand commonsense reasoning, where models must choose the most sensible continuation of a given scenario. PRobELM (Yuan et al., 2024b) assesses models’ capacity to rank scenarios by plausibility, bridging the gap between factual accuracy and world knowledge. While these benchmarks provide insights into models’ reasoning abilities, they do not address the challenges of forecasting future events.
@@ 7 Conclusion
We introduce FOReCAst, a benchmark for evaluating both forecasting accuracy and confidence calibration in language models. Unlike existing datasets, FOReCAst explicitly assesses confidence alongside predictions. Our results show that current models struggle with both prediction and well-calibrated confidence, underscoring the need for improved uncertainty estimation and confidence calibration.
Limitations
While FOReCAst represents a significant step toward evaluating forecasting accuracy and confidence calibration, there are inherent limitations that we view as opportunities for future research rather than fundamental shortcomings of our work. First, our dataset is constructed solely from Metaculus, which may not fully represent the global diversity of forecasting practices or question domains. Second, our method for deriving gold confidence relies on community forecasts and heuristic transformations that might not capture all nuances of human uncertainty. Lastly, our focus on English-language forecasts limits the benchmark’s applicability across different languages and cultural contexts. Addressing these issues is part of our future work agenda.
Ethical Statement
FOReCAst is built from data sourced exclusively from Metaculus, an English-language forecasting platform. As a result, the dataset may embody the linguistic, cultural, and social biases inherent in its user community. These biases could affect both question selection and confidence judgments. We acknowledge these concerns and stress that our benchmark is intended as an initial step toward more inclusive forecasting evaluations. Future efforts should aim to incorporate data from a broader range of platforms and languages to mitigate these biases.
@@ D.2 Evaluation Hyperparameters
The scaling factor $\alpha$ in Equation 4, Equation 8, and Equation 9 is set to 0.05. For Equation 5 and Equation 6, we set $\sigma_{\max}$ to 30 and $\sigma_{\min}$ to 1 for Timeframe Prediction, while for Quantity Estimation, $\sigma_{\max}$ is 20 and $\sigma_{\min}$ is 1. These values ensure that evaluation metrics appropriately scale errors and confidence calibration.

### v3

@@ 6 Related Work
Recent forecasting benchmarks focus on event prediction but largely overlook confidence calibration. OpenForecast (Wang et al., 2025) introduces a large-scale dataset for open-ended, multi-step event forecasting but does not assess model confidence or calibration. ForecastBench (Karger et al., 2024) evaluates binary (Yes/No) forecasting by prompting models for direct probability estimates, but since it queries each option independently, the assigned probabilities do not necessarily sum to 1, leading to potential inconsistencies. Both benchmarks advance the study of future event prediction but stop short of systematically evaluating uncertainty quantification, a crucial aspect for reliable deployment in decision-making contexts. Our work complements these efforts by providing a structured setting for evaluating both forecasting accuracy and probabilistic calibration.
Beyond forecasting, several benchmarks assess language models’ reasoning and inference capabilities. COPA (Roemmele et al., 2011) evaluates causal reasoning by presenting a premise and two alternatives, requiring models to select the more plausible cause or effect. HellaSwag (Zellers et al., 2019) challenges models with sentence completion tasks that demand commonsense reasoning, where models must choose the most sensible continuation of a given scenario. PRobELM (Yuan et al., 2024b) assesses models’ capacity to rank scenarios by plausibility, bridging the gap between factual accuracy and world knowledge. While these benchmarks probe useful reasoning skills, they are not designed for evaluating predictions about unresolved or future outcomes, nor do they assess confidence calibration in a dynamic context.
7 Limitations
FOReCAst is constructed entirely from Metaculus, a platform known for its high-quality, well-resolved forecasting questions and active user base. This makes it a strong and consistent foundation for developing the task of probabilistic forecasting with LLMs. However, it does not cover the full diversity of forecasting practices, such as different question formats, domain emphases, or forecasting styles seen in other platforms like Good Judgment Open or domain-specific expert settings. In this work, our primary focus is to define the task rigorously and establish a benchmark that is methodologically clean and scalable. Expanding to additional data sources and platforms is part of our future work and should help test generalization across forecasting cultures and environments.
Our method for constructing gold confidence labels is based on community-aggregated forecasts, which offer a practical and scalable proxy for crowd belief and uncertainty. While this approach enables systematic evaluation across hundreds of questions, it may not fully capture the variability and richness of individual forecasters’ reasoning or levels of disagreement. Moreover, the aggregated predictions may sometimes be influenced by group dynamics or correlated sources of error. We acknowledge these trade-offs, and view our current approach as a reasonable starting point given the limited availability of high-quality, large-scale confidence annotations. Future work could explore incorporating richer uncertainty representations, such as forecast distributions, forecaster rationales, or disagreement metrics.
Finally, our dataset is currently limited to English-language questions and forecasts, reflecting the language used by Metaculus and many existing NLP resources. While this is a reasonable initial scope for establishing the task, we recognize that forecasting behavior, uncertainty expression, and interpretation of probabilities may vary significantly across linguistic and cultural contexts. Extending FOReCAst to include multilingual questions and community forecasts is a promising direction for improving the inclusiveness and robustness of future benchmarks. We also note that some dimensions, such as temporal adaptation and real-time reasoning, remain out of scope here but are essential long-term goals for building forecasting-capable systems.
@@ 8 Conclusion
We introduce FOReCAst, a benchmark for evaluating both forecasting accuracy and confidence calibration in language models. Unlike existing datasets, FOReCAst explicitly assesses confidence alongside predictions. Our results show that current models struggle with both prediction and well-calibrated confidence, underscoring the need for improved uncertainty estimation and confidence calibration. Our dataset is now publicly available.
@@ Appendix B Dataset Statistics
Table 9: Top 10 months and categories across the full dataset and each question type. Note that questions can belong to multiple categories, reflecting the diversity of domains represented.
Global Distribution Timeframe Prediction
Month Count % Month Count %
2023-12 187 8.29% 2021-05 9 6.47%
2024-12 105 4.65% 2021-03 7 5.04%
2022-12 89 3.95% 2022-12 6 4.32%
2021-01 65 2.88% 2024-01 6 4.32%
2024-01 62 2.75% 2022-02 6 4.32%
2024-05 56 2.48% 2021-04 5 3.60%
2025-01 49 2.17% 2024-05 4 2.88%
2023-01 48 2.13% 2022-09 4 2.88%
2020-11 47 2.08% 2021-12 4 2.88%
2020-03 41 1.82% 2021-01 4 2.88%
Category Count % Category Count %
Politics 571 17.03% Health & Pandemics 40 18.10%
Geopolitics 486 14.50% Economy & Business 31 14.03%
Economy & Business 476 14.20% Politics 26 11.76%
Health & Pandemics 313 9.34% Technology 23 10.41%
Technology 253 7.55% Computing and Math 21 9.50%
Elections 218 6.50% Geopolitics 19 8.60%
Computing and Math 192 5.73% Artificial Intelligence 17 7.69%
Artificial Intelligence 151 4.50% Space 13 5.88%
Natural Sciences 139 4.15% Law 7 3.17%
Law 104 3.10% Natural Sciences 6 2.71%
Quantity Estimation Boolean Questions
Month Count % Month Count %
2021-02 25 6.96% 2023-12 167 9.50%
2022-12 21 5.85% 2024-12 94 5.35%
2020-04 19 5.29% 2022-12 62 3.53%
2023-12 18 5.01% 2024-01 51 2.90%
2024-05 17 4.74% 2021-01 45 2.56%
2020-03 17 4.74% 2025-01 44 2.50%
2021-01 16 4.46% 2023-01 38 2.16%
2020-11 16 4.46% 2024-05 35 1.99%
2021-10 12 3.34% 2023-06 34 1.93%
2020-12 12 3.34% 2024-11 29 1.65%
Category Count % Category Count %
Economy & Business 102 22.57% Politics 488 18.22%
Health & Pandemics 88 19.47% Geopolitics 428 15.98%
Politics 57 12.61% Economy & Business 343 12.80%
Geopolitics 39 8.63% Technology 205 7.65%
Artificial Intelligence 27 5.97% Elections 190 7.09%
Elections 25 5.53% Health & Pandemics 185 6.91%
Technology 25 5.53% Computing and Math 151 5.64%
Computing and Math 20 4.42% Natural Sciences 124 4.63%
Environment & Climate 16 3.54% Artificial Intelligence 107 3.99%
Sports & Entertainment 13 2.88% Law 96 3.58%
@@
Table 9 provides the top 10 categories and closing months for each task type, illustrating the breadth of topics and temporal diversity in FOReCAst.
@@ D.2 Evaluation Hyperparameters
The scaling factor $\alpha$ in Equation 3 and Equation 6 is set to 0.05. For Equation 4, we set $\sigma_{\max}$ to 30 and $\sigma_{\min}$ to 1 for Timeframe Prediction, while for Quantity Estimation, $\sigma_{\max}$ is 20 and $\sigma_{\min}$ is 1. These values ensure that evaluation metrics appropriately scale errors and confidence calibration.